### v1

# PAC-Bayes Analysis for Recalibration in Classification

Abstract
Nonparametric estimation with binning is widely employed in the calibration error evaluation and the recalibration of machine learning models. Recently, theoretical analyses of the bias induced by this estimation approach have been actively pursued; however, the understanding of the generalization of the calibration error to unknown data remains limited. In addition, although many recalibration algorithms have been proposed, their generalization performance lacks theoretical guarantees. To address this problem, we conduct a generalization analysis of the calibration error under the probably approximately correct (PAC) Bayes framework. This approach enables us to derive a first optimizable upper bound for the generalization error in the calibration context. We then propose a generalization-aware recalibration algorithm based on our generalization theory. Numerical experiments show that our algorithm improves the Gaussian-process-based recalibration performance on various benchmark datasets and models.
1 Introduction
Increasing the reliability of machine learning models is crucial in risk-sensitive applications such as autonomous driving [4]. Recently, the concept of calibration has become a significant measure of reliability, especially in classification tasks. In this context, the calibration performance is evaluated by how well predictive probabilities provided by our model align with the actual frequency of true labels. A close correspondence between them indicates that the model is well calibrated [6, 33]. To evaluate the calibration performance, the calibration error (CE) [11, 27, 9] or the top-label CE (TCE) [18, 9] is often used. This evaluates the disparity between the predicted probability of a model and the conditional probability of the label frequency given by the model prediction. Analytically computing the CE or TCE is, however, challenging because the conditional probabilities of the labels are intractable. A common approach to addressing this issue is constructing its estimator called the expected calibration error (ECE) by binning [35].
If the ECE is small, we consider the model well calibrated and its predictions highly reliable. Conversely, if the model is poorly calibrated, we correct the predictive probability by a post processing method called recalibration [10, 35] using a recalibration dataset that is independent of the train and test datasets. In the recalibration context, it is significant to theoretically guarantee that the recalibrated model can indeed achieve good calibration in terms of the ECE for unknown data. However, no studies have been conducted from this perspective to the best of our knowledge, although there has been extensive research into examining the estimation bias of the ECE from the CE [18, 11, 9, 29] and proposing specific algorithms for recalibration [24, 36, 23, 17, 10, 32, 38]. Additionally, no research has been conducted to analyze the generalization nature of the ECE in the case of nonrecalibration. This implies that there is no theoretical guarantee that a model achieving a small ECE on training data also achieves a small ECE on unknown data. In other words, regardless of recalibration, the generalization theory under the ECE is not yet established.
In this study, we construct a novel generalization theory for the ECE before and after recalibration using probably approximately correct (PAC) Bayes theory [21, 1], which is the common method of deriving generalization bounds across a wide range of models. Since PAC-Bayes theory can analyze generalization performance in an algorithm-dependent manner, the obtained upper bound is often optimizable and useful for deriving new generalization-aware algorithms. In applying PAC-Bayes analysis to the ECE, we face the following two challenges: (i) Owing to the nonparametric estimation of conditional probabilities, the ECE computed on the test dataset is not a sum of independent and identically distributed (i.i.d.) random variables and thus cannot be simply applied using the existing PAC-Bayes bounds derived under the i.i.d. assumption, and (ii) some bins may not contain samples because binning divides the probability interval $[0,1]$ into equal widths, making it difficult to apply the concentration inequality used in the PAC-Bayes bound derivation.
Our main contribution is the establishment of a generalization error bound in the ECE, including recalibration scenarios, by deriving new concentration inequalities that address the above challenges. This theory also enables the analysis of the bias between the ECE and the CE or the TCE through PAC-Bayes analysis. The results reveal the convergence rate of the generalization error and bias, $\mathcal{O}(1/n^{1/3})$, regardless of whether recalibration is performed, where $n$ is the number of data points for training or recalibration. Furthermore, our theory makes it possible to analyze the ECE in both binary and multiclass classifications. We then show that although the ECE can only capture the calibration performance for the top-$1$ predictive probability, it is unaffected by the curse of dimensionality.
Furthermore, we formulate a novel recalibration algorithm by leveraging our PAC-Bayes bound, which is expected to improve the generalization performance in terms of ECE and reduce bias in CE or TCE. Our method can be regarded as an extension of the recently proposed recalibration method based on Gaussian processes [32], incorporating flexible a Kullback–Leibler (KL) regularization with control parameters and calibration errors as a recalibration loss. Numerical experiments confirm a correlation between the KL terms in our bounds and the generalization performance, show that our method can improve GP-based recalibration, and reveal the instability of ECE-based calibration performance evaluation due to the slow convergence rate of $\mathcal{O}(1/n^{1/3})$.
2 Preliminaries
In this section, we summarize the basic notations, problem setup (Section 2.1), calibration metric (Section 2.2), and the generalization notion in recalibration (Section 2.3).
2.1 Notations and problem setting
For a random variable denoted in capital letters, we express its realization with corresponding lowercase letters. Let $P(X)$ denote a probability of $X$, and let $P(Y|X)$ represent the conditional probability of $Y$ given $X$. We express the expectation of a random variable $X$ as $\mathbb{E}_{X}$. Let $\mathrm{KL}(P\|Q)$ be the Kullback–Leibler (KL) divergence of $P$ from $Q$, where $Q$ is a probability distribution and $P$ is absolutely continuous with respect to $Q$.
Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ be the domain of data, where $\mathcal{X}$ and $\mathcal{Y}$ are input and label spaces, respectively. Suppose $\mathcal{D}$ represents an unknown data distribution, and let $S_{\mathrm{tr}}\coloneqq\{Z_{m}\}_{m=1}^{n_{\mathrm{tr}}}$ denote the training dataset consisting of $n_{\mathrm{tr}}$ samples drawn independently and identically distributed (i.i.d.) drawn from $\mathcal{D}$. We also define the test dataset comprising $n_{\mathrm{te}}$ samples as $S_{\mathrm{te}}\sim\mathcal{D}^{n_{\mathrm{te}}}$. Let $f_{w}:\mathcal{X}\to\mathcal{Y}$ be our parametric predictor and we denote its parameters as $w\in\mathcal{W}\subset\mathbb{R}^{d}$. We evaluate the performance of the trained predictor $f_{w}$ using the loss function $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, where $l(y,f_{w}(x))$ denotes the loss incurred by the prediction $f_{w}(x)$ for the target $y$. We define the training loss as $\hat{L}(w,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}l(Y,f_{w}(X))$ and the (expected) test loss as $L(w,\mathcal{D})\coloneqq\mathbb{E}_{S_{\mathrm{te}}}\frac{1}{n_{\mathrm{te}}}\sum_{(X,Y)\in S_{\mathrm{te}}}l(Y,f_{w}(X))$. Under this setting, the generalization error for $l$ is defined as $\mathrm{gen}(f_{w},S_{\mathrm{tr}},l)\coloneqq L(w,\mathcal{D})-\hat{L}(w,S_{\mathrm{tr}})$. The following theorem is the most basic form of the PAC-Bayes generalization bounds [1].
Theorem 1.
For any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and for any $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$, we have
$\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},l)\leq\frac{\mathrm{KL}(\rho\|\pi)+\log\mathbb{E}_{\pi,S_{\mathrm{tr}}}e^{\lambda(L(w,\mathcal{D})-\hat{L}(w,S_{\mathrm{tr}}))}+\log\frac{1}{\epsilon}}{\lambda}.$ (1)
2.2 Calibration metrics
Calibration for binary classification:
Let $\mathcal{Y}=\{0,1\}$ and let $f_{w}:\mathcal{X}\to[0,1]$ be a model, where the output corresponds to the model’s confidence that the label is $1$. In the context of calibration, it is expected that not only $f_{w}$ has high accuracy, but also its predictive probability aligns well in the actual label frequency $P(Y|f_{w}(X))$. A model is well calibrated [33, 12] when $P(Y|f_{w}(X))=f_{w}(X)$ holds; however, it is difficult to confirm this. Therefore, an alternative metric called as the calibration error (CE) [33, 12] is typically used. Given $W=w$, the CE is defined as
$\mathrm{CE}(f_{w})\coloneqq\mathbb{E}|P(Y=1|f_{w}(X))-f_{w}(X)|=\mathbb{E}|\mathbb{E}[Y|f_{w}(X)]-f_{w}(X)|.$ (2)
Unfortunately, evaluating the CE is still infeasible because $\mathbb{E}[Y|f_{w}(X)]$ is intractable. To resolve this issue, we evaluate the calibration performance by constructing the estimator of the CE. The widely used estimator is the expected calibration error (ECE), where binning is used to estimate $\mathbb{E}[Y|f_{w}(X)]$ [10, 35, 36]. The ECE is calculated by partitioning the range of the predictive probability $[0,1]$ into $B$ intervals =$\mathcal{I}=\{I_{i}\}_{i=1}^{B}$ (called bins) and averaging within each bin using an evaluation dataset $S_{e}\coloneqq\{z_{m}\}_{m=1}^{n_{e}}\in\mathcal{Z}^{n_{e}}$, where we assume $n_{e}\geq 2B$. That is, the ECE is defined as
$\mathrm{ECE}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}|\bar{f}_{i,S_{e}}-\bar{y}_{i,S_{e}}|,$ (3)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}$, $p_{i}\coloneqq\frac{|I_{i}|}{n_{e}}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}f_{w}(x_{m})$, and $\bar{y}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}y_{m}$. In the above, bins are set by dividing the interval $[0,1]$ into $B$ bins of the same width: $I_{1}=(0,1/B],I_{2}=(1/B,2/B],\dots,I_{B}=((B-1)/B,B]$. Setting $S_{e}=S_{\mathrm{te}}$, for instance, corresponds to evaluating the ECE on the test dataset.
Calibration for multiclass classification:
Let the label space be $Y\in\mathcal{Y}\coloneqq\{0,\dots,K-1\}$, where $K\in\mathbb{N}$ is the number of classes. We then define the one-hot encoding of the label and its domain as $\hat{Y}\in\hat{\mathcal{Y}}\subset\mathbb{R}^{K}$. Let $f_{w}:\mathcal{X}\to\Delta^{K}$ be a probabilistic classifier, where $\Delta^{K}$ represents the $K-1$ dimensional simplex. For example, such a simplex is obtained by the final softmax layer in neural networks. Under these settings, $f_{w}$ predicts the label by $C\coloneqq\mathrm{argmax}_{k}f_{w}(X)_{k}$, where $f_{w}(X)_{k}$ is the $k$-th dimension of $f_{w}(x)$, which represents the model’s confidence of the label $k\in K$.
A frequently used measure in the above setup is the top-label calibration error (TCE) [18, 9], which is defined as follow with respect to the highest prediction probability output by $f_{w}$:
$\mathrm{TCE}(f_{w})\coloneqq\mathbb{E}|P(Y=C|f_{w}(X)_{C})-f_{w}(X)_{C}|.$ (4)
Similarly to $\mathrm{CE}(f_{w})$, the estimator of the TCE can be obtained by binning to $f_{w}(X)_{C}$. That is, we divide the range of the predictive probability of $f_{w}(X)_{C}$ into $B$ bins of the same width, and then calculate the following ECE: $\mathrm{ECE}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|$, where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $p_{i}\coloneqq|I_{i}|/n_{e}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. For brevity, we refer to the TCE estimator in both binary and multiclass classifications constructed by binning as the ECE.
2.3 Generalization for recalibration analysis
The trained predictor $f_{w}(x)$ can be poorly calibrated [10, 18]. One common approach to address this problem involves recalibrating $f_{w}(x)$ by post processing using the parametric function  [^1] In the binary case, given $x$, we use $(f_{w}(x),1-f_{w}(x))\in\Delta^{2}$ as the input to $\eta_{V}(\cdot)$; see Appendix B for details. $\eta_{V}:\Delta^{K}\to\Delta^{K}$, where $V\in\mathcal{V}\subset\mathbb{R}^{d^{\prime}}$ is the parameter learned by the recalibration dataset $S_{\mathrm{re}}\sim\mathcal{D}^{n_{\mathrm{re}}}$ at fixed $w$ and $S_{\mathrm{re}}$ is independent of $S_{\mathrm{tr}}$. In this procedure, the overall dataset is split into training data $S_{\mathrm{tr}}$ for learning $W$, the recalibration data $S_{\mathrm{re}}$ for learning $V$, and the test data $S_{\mathrm{te}}$ for the evaluation of the ECE. We define the recalibrated model as $\eta_{v}(f_{w})$.
We expect $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})$ to be sufficiently small after recalibration. Therefore, it is important to theoretically explore the conditions for achieving an equally small $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})$ for the unknown test data $S_{\mathrm{te}}$ when $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})$ is sufficiently small. To this end, we study the following error:
$\mathrm{gen}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{ECE})\coloneqq|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|.$ (5)
We refer to this quantity as the generalization error of the ECE under $S_{\mathrm{re}}$.
Given that the ECE is an estimator of the CE or the TCE obtained by binning, it is important to evaluate how recalibration affects the bias between them expressed as
$\mathrm{Bias}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\coloneqq|\mathrm{TCE}(\eta_{v}(f_{w}))-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|.$ (6)
We refer to this as the estimation bias of the ECE under $S_{\mathrm{re}}$. Similarly, we define the estimation bias of CE as $\mathrm{Bias}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{CE})\coloneqq|\mathrm{CE}(\eta_{v}(f_{w}))-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|$ for binary classification.
3 PAC-Bayes bounds for ECE without recalibration
Our goal is to provide a comprehensive analysis of Eqs. (5) and (6) by utilizing the PAC-Bayes theory. Before showing our main results for them, we present our PAC-Bayes analysis for the following two quantities: $\mathrm{gen}(f_{w},S_{\mathrm{tr}},\mathrm{ECE})\coloneqq|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ and $\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})\coloneqq|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$, for which we do not conduct recalibration. The results of this analysis are crucial for our recalibration analysis in Section 4 and provide a rigorous understanding of the generalization and estimation bias in the ECE in the absence of recalibration.
3.1 Generalization error analysis for calibration
We show the first PAC-Bayes bound for the generalization error of the ECE as follows.
Theorem 2.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. For both binary and multiclass settings, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (7)
Proof sketch.
We describe the case for binary classification. We first derive the following reformulation: $\mathrm{ECE}(f_{w},S_{\mathrm{tr}})=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|$, where $\mathbb{E}_{\widehat{S}_{\mathrm{tr}}}$ is the empirical expectation for $S_{\mathrm{tr}}$. We can derive a similar equation for $S_{\mathrm{te}}$. By using the triangular inequality and setting $l_{i}(Z)=(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}$, we have
$|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)-\mathbb{E}_{Z\sim\widehat{S}_{\mathrm{tr}}}l_{i}(Z)\right|.$ (8)
By using McDiarmid’s inequality to evaluate the exponential moment of Eq. (8), we obtain
$\log\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}},\pi}e^{\lambda\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)-\mathbb{E}_{Z\sim\widehat{S}_{\mathrm{tr}}}l_{i}(Z)\right|}\leq B\log 2+4\lambda^{2}/n$ (9)
for any $\lambda\in\mathbb{R}$. Combining this result with the Donsker–Varadhan inequality leads to Eq. (7). ∎
The complete proof is provided in Appendix A.2. This theorem shows that the generalization properties in the ECE behave similarly for the binary and multiclass cases. This is natural, given that the binning-based ECE divides the top-$1$ predicted probabilities equally into $B$ bins in both cases.
Discussion about the assumption of $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$: Although a single data point can be used to evaluate loss in a standard generalization error analysis, multiple data points are required to evaluate the ECE, as it is an estimator. Therefore, for a fair comparison, the same number of data points should be used in the analysis of the ECEs based on $S_{\mathrm{tr}}$ and $S_{\mathrm{te}}$. In this case, it seems difficult to apply the common method of deriving tighter PAC-Bayes bounds, which involves analyzing the KL divergence of the Bernoulli random variable $kl$ [20, 7]. This is because the ECE at $S_{\mathrm{te}}$ is no longer a sum of i.i.d. random variables as it is in a test loss.
Comparison with existing work: The most closely related analysis is provided by Gupta and Ramdas [11]. By comparison, our analysis provides a more general and tighter bound, as explained below. First, they analyzed the ECE for binary classification under the setting where an equal number of samples are allocated in each bin. In contrast, our analysis covers both binary and multiclass classifications under the equal-width binning setting, which is more common in practice. Moreover, their analysis does not include the generalization error analysis, but they rather derived the following concentration bound of the ECE: $|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{val}})|\leq\mathcal{O}_{p}\left(\sqrt{B\log B/n}\right)$, which corresponds to the case where $S_{e}=S_{\mathrm{val}}\sim\mathcal{D}^{n}$ as in our study. Furthermore, when setting $\lambda=\sqrt{BN}$, our bound exhibits $\mathcal{O}_{p}(\sqrt{B/n})$ that improves the order with $\log B$. IN their proof strategy, they evaluate $y-f_{w}(x)$ for each bin, necessitating multiple applications of concentration inequalities. Therefore, aggregating these errors across all bins using a union bound results in a slow convergence. When we adopt this proof technique in our analysis, the resulting bound exhibits $\mathcal{O}_{p}(\sqrt{B\times\mathrm{KL}(\rho\|\pi)/n})$, which is significantly larger than the order of Eq. (7). In comparison, our bound simultaneously treats all bins using Eq. (8) and we should only use Eq. (9) once, resulting in an improved order.
3.2 Estimation bias analysis for calibration
Next, we show the result of our estimation bias analysis under the following smoothness assumption, which is commonly used in the nonparametric estimation [30].
Assumption 1.
Conditioned on $W=w$, the $L$-Lipschitz continuity holds for $\mathbb{E}[Y|f_{w}(x)]$ in the binary classification and $\mathbb{E}[Y|f_{w}(x)_{C}]$ in the multiclass classification.
Theorem 3.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{W},S_{\mathrm{tr}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (10)
Exactly the same statement holds for $\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})$ (see Appendix A.3 for the statement).
Proof sketch.
To analyze the bias caused by binning, following Kumar et al. [18], we define the conditional expectation of $f_{w}$ given bins as $f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}$. With this definition, we decompose the estimation bias as follows: $|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}}^{C})|+|\mathrm{TCE}(f_{\mathcal{I}}^{C})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$. We then evaluate these terms separately. The first term on the right-hand side is the bias caused by binning, which can be bounded using Assumption 1. The second term can be upper-bounded by $\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})$ using Jensen’s inequality. See Appendix A.3 for details. ∎
According to this theorem, the number of bins shows a trade-off relationship as follows. As we increase $B$, the bias caused by binning decreases because we estimate the label frequency more precisely. On the other hand, $\mathbb{E}_{\rho}\mathrm{gen}(f_{W},S_{\mathrm{tr}},\mathrm{ECE})$ increases. From this observation, we derive the optimal number of bins that minimizes the estimation bias. Assuming that $\mathrm{KL}(\rho\|\pi)$ is sufficiently smaller than $n$, for example, $\mathcal{O}(\log n)$ and independent of $B$, and setting $\lambda=\sqrt{BN}$, we can minimize the upper bound of Eq. (10) with respect to $B$. This yields $B=\mathcal{O}(n^{1/3})$, and under this bin size, the order of the estimation bias becomes
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{W})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|=\mathcal{O}(\log n/n^{1/3}).$ (11)
Similarly, we can also obtain the same optimal bin size for binary classification.
Relationship with nonparametric regression: The order of the estimation bias in Eq. (11) is tight from the the nonparametric regression viewpoint. Recall that the TCE measures the error between two function: $\mathbb{E}[Y|f_{w}(x)]$ and $f_{w}(x)$. We estimate $\mathbb{E}[Y|f_{w}(x)]$ by binning, which is a nonparametric method. Thus, this is a problem of nonparametric regression on $[0,1]$ under Lipschitz continuity. According to Tsybakov [30], the error in nonparametric regression cannot be smaller than $\mathcal{O}(1/n^{1/3})$. Achieving an order smaller than this necessitates additional assumptions about the underlying data distribution. In this regard, the order of our bound is convincing under the current assumptions. We note that assuming Hölder continuity instead of Assumption 1 does not improve the order of the estimation bias owing to the bias caused by binning (see Appendix A.4 for details).
3.3 Discussion about top-$K$ calibration
In applications such as medical diagnosis [14], it is often important to calibrate the prediction probabilities for any $k\in K$, not simply focusing on the top-$1$. In this case, the alternative metric $CE_{K}\coloneqq\mathbb{E}\|P(\hat{Y}|f_{w}(X))-f_{w}(X)\|_{1}$ has been explored instead of TCE [9], where $\|\cdot\|_{p}$ is the $L^{p}$ distance in $\mathbb{R}^{K}$. The $CE_{K}$ metric measures the calibration performance of models across all classes. Here, we show that the direct estimation of $CE_{K}$ by binning results in a slow convergence due to the curse of dimensionality.
Let us consider the following binning scheme to estimate $CE_{K}$. We split each dimension $[0,1]$ into $B^{\prime}$ bins of the same width because we want to estimate the $K$-dimensional conditional probability in $[0,1]^{K}$ by binning. After doing so, there are $B=(B^{\prime})^{K}$ bins, and $[0,1]^{K}$ is split into $B$ small regions. We refer to these regions as $\mathcal{I}\coloneqq\{I_{i}\}_{i=1}^{B}$. In this case, the ECE of $CE_{K}$ can be expressed as $\mathrm{ECE}_{K}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}\frac{\#B_{i}}{n_{e}}\|\bar{f}_{i,S_{e}}-\bar{y}_{i,S_{e}}\|_{1}$, where #$\#B_{i}$ is the number of instances in $I_{i}$, $\bar{f}_{i,S_{e}}$ is the empirical mean of the prediction, and $\bar{y}_{i,S_{e}}$ is the empirical mean of the one-hot encoding of the label in $I_{i}$. Under these settings, the upper bound of the estimation bias in $\mathrm{ECE}_{K}$ is given as follows.
Theorem 4.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$ and $f_{w}(x)$ has the probability density over $[0,1]^{K}$ and $\mathbb{E}[\hat{Y}|f_{w}(x)]$ satisfies $L$-Lipschitz continuity. Then, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|CE_{K}(f_{W})-\mathrm{ECE}_{K}(f_{W},S_{\mathrm{tr}})|\leq\frac{K(1+L)}{B^{\frac{1}{K}}}+\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{K^{2}\lambda^{2}}{2n}}{\lambda}.$ (12)
By minimizing the upper bound of Eq. (12), we obtain the optimal bin size as $B=\mathcal{O}(n^{\frac{1}{K+2}})$, leading to an estimation bias of $\mathcal{O}(K^{\frac{3}{2}}\log n/n^{\frac{1}{K+2}})$. Although this order aligns with the lower bound of $K$-dimensional nonparametric regression with Lipschitz continuity [30], it is significantly larger than that of the TCE due to the curse of dimensionality. On the other hand, the TCE successfully circumvents this issue as described in Theorem 3 by focusing only on the top-$1$ predicted label, although it does not necessarily guarantee good calibration performance for all $K$ [9].
4 PAC-Bayes bounds under recalibration
In this section, we extend the bounds from Section 3.1 to the recalibration setting, leading to our novel recalibration algorithm in Section 4.2. All the proofs are provided in Appendix B.
4.1 Bounds for ECE and estimation bias under recalibration
The following result shows the PAC-Bayes bound of the ECE under the recalibration context.
Corollary 1.
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$. For both binary and multiclass classifications, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior $\tilde{\rho}$ over $V$, and for any $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{ECE})\leq\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}$ (13)
This result implies that the KL regularization over the parameter space is required to prevent overfitting similarly to the PAC-Bayes bound when using the training dataset [21, 1]. Since the recalibration data is available and we have fixed $w$, we only consider the posterior $\tilde{\rho}$ over $V$. Instead of fixing $W=w$, we can also obtain the PAC-Bayes bound by taking expectation over $W$ (see Appendix B.2).
Comparison with Theorem 2: $\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})$ is usually smaller than $\mathrm{KL}(\rho\|\pi)$ because of the following two reasons. First, the dimension $d^{\prime}$ of $V$ is much smaller than $d$. This is because, for example, $d^{\prime}$ typically coincides with the number of labels $K$ in the multiclass classification, whereas $d$ is the number of parameters $W$ in large models such as DenseNet [13] used in Section 6. Second, the prior $\tilde{\pi}$ can be dependent on $S_{\mathrm{tr}}$ since $S_{\mathrm{re}}$ and $S_{\mathrm{tr}}$ are independent, which means that we can use a data-dependent prior. These facts suggest that recalibration can improve generalization in terms of the ECE.
Next, we present our PAC-Bayes bound on the estimation bias of the ECE and TCE in the recalibration context after introducing the necessary assumption similar to Assumption 1.
Assumption 2.
Conditioned on $W=w$ and $V=v$, the $L$-Lipschitz continuity holds for $\mathbb{E}[Y|\eta_{v}(f_{w}(X))]$ in binary classification and $\mathbb{E}[Y|\eta_{v}(f_{w}(X))_{C}]$ in the multiclass classification.
Corollary 2.
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$. Under Assumption 2, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$, and for any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}$ (14)
Exactly the same statement holds for $\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{CE})$ (see Appendix B.1).
Similarly to Eq. (11), there is a trade-off relationship with respect to $B$ in the above bound. By minimizing this upper bound with respect to $B$, we obtain the optimal bin size as $B=\mathcal{O}(n^{1/3})$. For practical purposes, it is possible to set $B=\lfloor n^{1/3}\rfloor$, where $\lfloor x\rfloor\coloneqq\max\{m\in\mathbb{Z}:m\leq x\}$.
4.2 Proposed recalibration algorithm based on PAC-Bayes bounds
[…]
However, minimizing this function is difficult since the ECE is neither smooth nor differentiable. To avoid this issue, we focus on the following fact: $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}\|\hat{Y}-\eta_{v}(f_{w}(X))\|_{2}^{2}$ in the multiclass classification  [^2] A similar relation holds for binary classification. The complete proof can be found in Appendix B.3.. This upper bound the Brier score [9], which is a continuous and differentiable calibration metric. Therefore, under the posterior $\tilde{\rho}(v)=\tilde{\rho}(v;\theta)$ parametrized by $\theta$, we can minimize the following PAC-Bayes-based objective by gradient-based optimization:
$\min_{\theta}\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}\|\hat{Y}-\eta_{V}(f_{w}(X))\|^{2}_{2}+\alpha\frac{\mathrm{KL}(\tilde{\rho}(v;\theta)\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (15)
We expect the recalibrated models to achieve not only a low TCE or CE but also a high accuracy in practice. We thus consider minimizing the following objective with the classification loss $l_{\mathrm{acc}}$:
$\min_{\theta}\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}(l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2})+\alpha\frac{\mathrm{KL}(\tilde{\rho}(v;\theta)\|\tilde{\pi})}{n_{\mathrm{re}}}.$ (16)
Eq. (16) is also guaranteed through PAC-Bayes bounds in terms of the test accuracy and TCE (or CE) simultaneously because it can be derived by utilizing the union bound for Eq. (1) and Corollary 2. The detailed derivation can be seen in Corollary 5 in Appendix B.3.
We refer to the recalibration achieved by minimizing Eqs. (15) and (16) as PAC-Bayes recalibration (PBR) and summarize these overall procedures in Algorithm 1. As an example of a recalibration model in PBR, we can consider using the Gaussian process (GP) [25]. In this case, $\tilde{\rho}(v;\theta)$ is set as the multivariate Gaussian distribution, $\mathcal{N}(v;\mu,\Sigma)$, where $\mu$ and $\Sigma$ are the mean and covariance matrix, respectively. Furthermore, as discussed in Section 4.1, we can construct the GP prior $\tilde{\pi}$ using the output values of $f_{w}(x)$ for $x\in S_{\mathrm{re}}$ during training, where $w$ is learned using $S_{\mathrm{tr}}$. This is because such $\tilde{\pi}$ can be dependent on $S_{\mathrm{tr}}$ since $S_{\mathrm{tr}}$ and $S_{\mathrm{re}}$ are independent. In this way, we can construct a data-dependent prior, which is expected to yield a small KL value.
A similar GP-based recalibration model was also used by Wenger et al. [32]. In this study, we propose recalibration via variational inference (VI), i.e., minimizing the following objective function derived from the evidence lower bound (ELBO): $\mathbb{E}_{\tilde{\rho}(v;\theta)}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\mathrm{KL}(\tilde{\rho}(v;\theta)|\tilde{\pi})/n_{\mathrm{re}}$, where $l_{\mathrm{acc}}$ is defined as the softmax loss. Eq. (16) corresponds to this ELBO objective when $\alpha=1$ and the Brier score is removed. This fact indicates that PBR extends the GP-based method to have flexible KL regularization and Brier score loss to control calibration performance on the basis of the PAC-Bayes notion. In Section 6.2, we compare PBR and GP-based recalibration using the same recalibration model settings.
Algorithm 1 PAC-Bayes recalibration (PBR)
1:  INPUT: All dataset $S_{\mathrm{all}}$, model $f_{w}$, recalibration model $\eta_{v}$, variational posterior $\tilde{\rho}(v;\theta)$
2:  Splitting all dataset $S_{\mathrm{all}}$ to $S_{\mathrm{tr}}$, $S_{\mathrm{re}}$, $S_{\mathrm{te}}$.
3:  Training $f_{w}$ using $S_{\mathrm{tr}}$.
4:  Updating $\theta$ by minimizing Eq. (15) or (16) using $S_{\mathrm{re}}$ at fixed $w$.
5:  Obtaining $V$ by taking the mean of $J$ samples $\{V_{j}\}_{j=1}^{J}\overset{\mathrm{i.i.d.}}{\sim}\tilde{\rho}(v;\theta)$
6:  (Option) Calculate the test ECE: $\mathrm{ECE}(\eta_{V}(f_{w}),S_{\mathrm{te}})$ by setting $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$.
7:  RETURN: $\eta_{V}(f_{w})$, ($\mathrm{ECE}(\eta_{V}(f_{w}),S_{\mathrm{te}})$)
5 Related work
In classification, some exciting studies have utilized the PAC-Bayes theory to analyze various metrics other than accuracy. Ridgway et al. [26] studied the AUC metric. Morvant et al. [22] analyzed the confusion matrix in multiclass classification. We extend this line of inquiry to the classification metric by focusing on calibration. Additionally, in terms of uncertainty quantification, Sharma et al. [28] analyzed the conformal prediction, which is a different metric for uncertainty quantification.
Gupta and Ramdas [11] and Kumar et al. [18] examined the ECE with binning in binary classification focusing on uniform mass binning, which divides intervals of the same mass. They also analyzed a recalibration method based on this binning method. In contrast, our analysis of the ECE in a multiclass setting, which has not yet been performed, focuses on the typical equal-width binning method. Sun et al. [29] also analyzed the binning-based recalibration for binary classification under a similar Lipschitz assumption. They have shown the same estimation bias order as in our results only when the histogram binning recalibration is performed, suggesting $\mathcal{O}(1/n^{1/3})$. This result is expected from the viewpoint of nonparametric regression, as discussed in Section 3.1.
The analysis of multiclass calibration is limited compared with the binary classification. Zhang et al. [38] investigated the kernel-based $CE_{K}$. However, their analysis does not address generalization or recalibration, which fundamentally distinguishes our approach. Moreover, Zhang et al. [38] remarked that the binning-based $CE_{K}$ suffers from the curse of dimensionality without a formal statement, whereas our Theorem 4 validates this phenomenon. Gruber and Buettner [9] discussed various calibration metrics and clarified that the TCE is a weaker calibration metric than to $CE_{K}$. On the other hand, our Theorem 4 enables the non-asymptotic bias analysis, showing that the naive estimator of $CE_{K}$ suffers from the curse of dimensionality compared with the TCE.
6 Experiments
In this section, we confirm the correlation between our bounds in Corollary 1 and the generalization performance of recalibration (Section 6.1). In Section 6.2, we discuss the effectiveness of our PBR in Section 4.2. Owing to page limitation, we report here the results of the multiclass classification experiments on the MNIST [19] and CIFAR-100 [15] datasets. All experimental results can be found in Appendix D. In addition, the details of our experimental settings, including model information, are summarized in Appendix C. We also set $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$ for ECE evaluation on the basis of our findings.
6.1 Verification of our bounds
In this section, we empirically confirm the correlation between the KL divergence term in Eq. (13) and the generalization gap in the ECE. To achieve this, we conducted recalibration experiments using PBR presented in Eq. (15) with various constraints ($\alpha$) for the KL term. Then, we measured each KL value and ECE gap after recalibration by performing $10$-fold cross-validation, where we set $n_{\mathrm{re}}=1000$.
A portion of the results is shown in Figure 1. From these results, we can see that the KL and generalization gap of the ECE are somewhat correlated in terms of Pearson’s and Kendall’s rank-correlation coefficients, which validates our bound. It is also evident that the ECE gap is noisy. This is because the estimation bias of the ECE through binning has a very slow convergence rate of $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$, even under our optimal $B$. When using $n_{\mathrm{re}}=1000$, the noise due to the estimation bias can be $\mathcal{O}(1/10)$. This implies that the degree of correlation is likely to be buried in this noise if models achieve a small ECE of less than $0.1$. A simple way to reduce the noise is to increase the number of recalibration data points; however, this is impractical according to $\mathcal{O}(1/n_{\mathrm{re}}^{1/3})$ convergence rate.
6.2 Empirical comparison of recalibration methods
In this section, we compare our proposed PBR and existing standard or recent recalibration methods, temperature scaling [10] and GP calibration [32]. The optimizable parameter $\alpha$ in PBR is selected from $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 } by grid search.
We summarize the results in Table 1. These results show that our method effectively improves calibration performance over existing methods, especially the GP-based recalibration method. We also found that, contrary to the report by Wenger et al. [32], temperature scaling performs better than GP-based recalibration in some cases in terms of the ECE with the optimal $B$. The failure to reproduce this result could be because Wenger et al. [32] reported the ECE with $B=100$. According to our analysis, this setting achieves the smallest estimation bias when $n_{\mathrm{te}}=\mathcal{O}(100^{3})$. However, since the $n_{\mathrm{te}}$ actually used is far smaller than this (see Table 2 in Appendix C), severe bias may occur.
Despite these instabilities, PBR appears to consistently improve the existing GP methods, at least in the case of the ECE under the optimal $B$ studied in this paper. The results of other experiments shown in Appendix D confirmed that GP-based methods, including PBR, do not consistently outperform standard recalibration methods such as temperature scaling. This indicates the need to select a recalibration model appropriate for the dataset and model used.
Figure 1: KL vs ECE gap on MNIST (Multiclass; $1$-layer NN) and CIFAR-100 (Multiclass; ResNeXt-$29$ [34] and DenseNet-BC-190 [13]) dataset per various regularize parameters ($\alpha$).
Table 1: Results of multiclass classification experiments on MNIST and CIFAR-100. The symbols ($\downarrow$) and ($\uparrow$) refer to lower and higher values indicating better performance, respectively.
Data Model Uncalibrate GP [32] PBR (Ours; Eq. (15)) PBR (Ours; Eq. (16))) Temp. Scaling
ECE ($\downarrow$) Accuracy ($\uparrow$) ECE ($\downarrow$) Accuracy ($\uparrow$) ECE ($\downarrow$) Accuracy ($\uparrow$) ECE ($\downarrow$) Accuracy ($\uparrow$) ECE ($\downarrow$) Accuracy ($\uparrow$)
MNIST XGBoost .$0036\pm.0003$$9785\pm.0005$$0038\pm.0005$$9785\pm.0007$$\mathbf{.0035\pm.0005}$$9784\pm.0007$$0037\pm.0007$$9785\pm.0007$$0055\pm.0017$$9785\pm.0005$
random forest .$1428\pm.0007$$9659\pm.0005$$0324\pm.0036$$9663\pm.0011$$\mathbf{.0052\pm.0018}$$9660\pm.0009$$0066\pm.0029$$9661\pm.0012$$0062\pm.0009$$9659\pm.0005$
1layer NN .$0164\pm.0004$$9760\pm.0005$$01$0023 .$9740\pm.0008$$0137\pm.0014$$9759\pm.0007$$0112\pm.0019$$9740\pm.0016$$\mathbf{.0062\pm.0026}$$9760\pm.0005$
CIFAR100 alexnet .$2548\pm.0007$$4374\pm.0008$$2548\pm.0008$$4373\pm.0009$$2548\pm.0007$$4372\pm.0009$$2548\pm.0007$$4373\pm.0009$$\mathbf{.0216\pm.0037}$$4374\pm.0008$
WRN-28-10-drop .$0568\pm.0009$$8131\pm.0011$$0567\pm.0010$$8129\pm.0011$$0504\pm.0054$$8129\pm.0011$$\mathbf{.0349\pm.0093}$$8129\pm.0011$$0374\pm.0028$$8131\pm.0011$
resnext-8x64d .$0401\pm.0010$$8229\pm.0013$$0$0010 .$8229\pm.0014$$0320\pm.0050$$8228\pm.0014$$\mathbf{.0310\pm.0082}$$8230\pm.0014$$0401\pm.0019$$8229\pm.0013$
resnext-16x64d .$0405\pm.0012$$8231\pm.0013$$0403\pm.0013$$8230\pm.0014$$\mathbf{.0305\pm.0056}$$8231\pm.0013$$0321\pm.0071$$8230\pm.0014$$0420\pm.0025$$8231\pm.0013$
densenet-bc-L190-k40 .$0639\pm.0008$$8230\pm.0007$$0639\pm.0008$$8229\pm.0007$$0630\pm.0045$$8228\pm.0006$$0618\pm.0054$$8229\pm.0007$$\mathbf{.0222\pm.0013}$$8230\pm.0007$
7 Conclusion and limitations
This paper provides the first comprehensive analysis of the generalization error and estimation bias of the ECE in both binary and multiclass classifications, resulting in several non-asymptotic bounds for them and the practical optimal bin size. Moreover, we develop a new generalization-aware recalibration algorithm based on our PAC-Bayes bound. Our analysis also reveals the slow convergence of the ECE through binning, which is a fundamental limitation. As discussed in Section 3.1, the binning method cannot leverage the underlying smoothness of the data distribution. Our numerical results suggest that the limited number of test data points makes the numerical evaluation of the ECE unstable in practice. We believe that other nonparametric methods used to estimate the CE or TCE may resolve this issue. An investigation of this point is our important future work.
Acknowledgements
MF was supported by RIKEN Special Postdoctoral Researcher Program. MF was supported by JST, ACT-X Grant Number JPMJAX210K, Japan. FF was supported by JSPS KAKENHI Grant Number JP23K16948. FF was supported by JST, PRESTO Grant Number JPMJPR22C8, Japan.
Appendix A Proofs and discussion of Section 3
For the latter purpose, we define $[n]=\{1,\ldots,n\}$ for $n\in\mathbb{N}$ and $\lfloor x\rfloor=\max\{m\in\mathbb{Z}:m\leq x\}$.
A.1 Additional preliminary
We use the following lemma repeatedly in our proofs.
Lemma 1 (Used in the proof of McDiarmid’s inequality).
[2] We say that a function $f:\mathcal{X}\to\mathbb{R}$ has the bounded difference property if for some nonnegative constants $c_{1},\dots,c_{n}$,
$\sup_{x_{1},\dots,x_{n},x_{i}^{\prime}\in\mathcal{X}}|f(x_{1},\dots,x_{n})-f(x_{1},\dots,x_{i-1},x^{\prime}_{i},x_{i+1},\dots,x_{n})|\leq c_{i},\quad 1\leq i\leq n.$ (17)
If $X_{1},\dots,X_{n}$ are independent random variables taking values in $\mathcal{X}$ and $f$ has the bounded difference property with constants $c_{1},\dots,c_{n}$, then for any $t\in\mathbb{R}$, we have
$\mathbb{E}\left[e^{t(f(X_{1},\dots,X_{n})-\mathbb{E}[f(X_{1},\dots,X_{n})])}\right]\leq e^{\frac{t^{2}}{8}\sum_{i=1}^{n}c_{i}^{2}}.$ (18)
A.2 Proof of Theorem 2
A.2.1 Reformulations of the ECE to loss functions
Here, we introduce the reformulations of the ECE, which is the essential technique in our proof.
Binary classification
We first focus on the binary setting. The training ECE can be reformulated as the empirical mean of $Y-f_{w}(X)$ in each bin as follows
$\mathrm{ECE}(f_{w},S_{\mathrm{tr}})\coloneqq\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|,$ (19)
which follows immediately from the definitions. Here $\widehat{S}_{\mathrm{tr}}$ is the empirical distribution of the training dataset.
Recall the conditional expectation of $f_{w}$ given bins
$f_{\mathcal{I}}(x)\coloneqq=\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}.$ (20)
The following relation holds
$\mathrm{CE}(f_{\mathcal{I}})=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|,$ (21)
where $\mathrm{CE}(f_{\mathcal{I}})$ means the CE of the function $f_{\mathcal{I}}$. We can derive this as follows; By definition, we have
$\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|$
$=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\mathcal{D}}[(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}]|$
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}|\mathbb{E}[Y|f_{w}(X)\in I_{i}]-\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]|,$ (22)
where we used the definition of the conditional expectation. On the other hand, We have
$\mathrm{CE}(f_{\mathcal{I}})$ $=\mathbb{E}|\mathbb{E}[Y|f_{\mathcal{I}}(x)]-f_{\mathcal{I}}(x)|$
$=\sum_{i=1}^{B}\mathbb{E}\left[|\mathbb{E}[Y|f_{\mathcal{I}}(x)]-f_{\mathcal{I}}(x)|\cdot\mathbbm{1}_{f_{\mathcal{I}}(X)\in I_{i}}\right]$
$=\sum_{i=1}^{B}P(f_{\mathcal{I}}(x)\in I_{i})\mathbb{E}\left[|\mathbb{E}[Y|f_{\mathcal{I}}(x)]-f_{\mathcal{I}}(x)|f_{\mathcal{I}}(X)\in I_{i}\right]$
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}|\mathbb{E}[Y|f_{w}(X)\in I_{i}]-\mathbb{E}[f_{w}(X)\in I_{i}]|,$ (23)
where we used the tower property. This concludes the proof.
Thus, we can transform the loss and ECEs by Eqs. (19) and (21).
Multiclass classification
Next, we consider the ECE for the multiclass setting. Let $C\coloneqq\mathrm{argmax}_{k}f_{w}(X)_{k}$. Then by definition of ECE, we have
$\mathrm{ECE}(f_{w},S_{e})=\sum_{i=1}^{B}\frac{|I_{i}|}{n_{e}}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(\mathbbm{1}_{Y=C}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}},$ (24)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. As for the conditional function, we define
$f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}f_{I_{i}}^{C}(x)\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}=\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)_{C}\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}.$ (25)
Then by repeating the exactly same proof for Eq. (21), we obtain
$\mathrm{TCE}(f_{\mathcal{I}}^{C}(x))=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(\mathbbm{1}_{Y=C}-f_{w}(X)_{C})\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}|,$ (26)
We can see that the ECE of the binary and multiclass classification essentially represents the same quantity; in the multiclass setting, by re-labeling $C$ as $1$ and others 00, they represent the same quantity.
A.2.2 Preliminary of the proof of Theorem 2
Next, we introduce the assumption, which is used in existing analysis Gupta and Ramdas [11] and not necessarily in our proof;
Assumption 3.
Given $W=w$, $f_{w}(x)$ is absolutely continuous with respect to the Lebesgue measure.
This assumption means that $f_{w}(x)$ has a probability density, and $f_{w}(x)$, it is satisfied without loss of generality as elaborated in Appendix C in Gupta and Ramdas [11]. Although this assumption is not required for our main results, if we assume this, we can improve the coefficient of the upper bound. Thus in the below, we provide the proof with and without this assumption simultaneously.
Thanks to this assumption, $f_{w}(X)$ takes distinct values almost surely, so for example, the situation when all training samples take the same predicted probability will be circumvented. When considering the multiclass setting with $CE_{K}$, this property plays an important role in eliminating the curse of dimensionality, see Appendix A.6.
We start from the slightly different version of Theorem 2. Then we extend it  2.
Theorem 5.
Assume that $n_{\mathrm{tr}}=n$. For both binary and multiclass settings, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (27)
Furthermore, in addition to the above assumption, when Assumption 3 holds, we have
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{2n}}{\lambda}.$ (28)
Proof.
To simplify the notation, we express $S_{\mathrm{tr}}$ as $S$ in this proof.
We first prove the binary setting. To derive the PAC Bayes bound, we transform the loss using Eq. (19) and (21), we have
$|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|$
$=\left|\sum_{i=1}^{B}\left|\mathbb{E}_{Z^{\prime}=(X^{\prime},Y^{\prime})}\left[(Y^{\prime}-f_{W}(X^{\prime}))\cdot\mathbbm{1}_{f_{W}(X^{\prime})\in I_{i}}\right]\right|-\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{m=1}^{n}(Y_{m}-f_{W}(X_{m}))\cdot\mathbbm{1}_{f_{W}(X_{m})\in I_{i}}\right|\right|$
$\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z^{\prime}=(X^{\prime},Y^{\prime})}\left[(Y^{\prime}-f_{W}(X^{\prime}))\cdot\mathbbm{1}_{f_{W}(X^{\prime})\in I_{i}}\right]-\frac{1}{n}\sum_{m=1}^{n}(Y_{m}-f_{W}(X_{m}))\cdot\mathbbm{1}_{f_{W}(X_{m})\in I_{i}}\right|$
$\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m})\right|,$ (29)
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$.
We then evaluate the exponential moment as
$\mathbb{E}_{S,\pi}e^{t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|}\leq\mathbb{E}_{\pi}\mathbb{E}_{S}e^{t\sum_{i=1}^{B}\left|\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m})\right|}.$ (30)
By setting $g(i,S)\coloneqq\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m})$, we have
$\mathbb{E}_{S}e^{t\sum_{i=1}^{B}|g(i,S)|}$ $=\mathbb{E}_{S}\prod_{i=1}^{B}e^{t|g(i,S)|}$
$\leq\mathbb{E}_{S}\prod_{i=1}^{B}\left(e^{tg(i,S)}+e^{-tg(i,S)}\right)$
$\leq\mathbb{E}_{S}\sum_{v_{1},\dots,v_{B}=0,1}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}g(i,S)}$ (31)
$=\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}g(i,S)}$
$=\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m})\right]},$ (32)
where $\sum_{v_{1},\dots,v_{B}=0,1}$ is all the combinations that will be generated by expanding $\prod_{i=1}^{B}$ in Eq. (31) and it has $2^{B}$ combinations.
We would like to upper bound $\mathbb{E}_{S}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m})\right]}$ using Lemma 1. For that purpose, here we evaluate $c_{i}$s of Lemma 1. By focusing on the exponent, we can estimate $c_{i}$s by
$\sup_{\{z_{m}\}_{m=1},\tilde{z}_{m}\in\mathcal{Z}}\sum_{i=1}^{B}t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(z_{m}^{\prime})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m\neq m^{\prime}}^{n}l_{i}(z_{m}^{\prime})-\frac{1}{n}l_{i}(\tilde{z}_{m^{\prime}})\right]$
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\sum_{i=1}^{B}\frac{t(-1)^{v_{i}}}{n}\cdot\left[-l_{i}(z^{\prime}_{m^{\prime}})+-l_{i}(\tilde{z}_{m^{\prime}})\right]$
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\frac{t(-1)^{v_{1}}}{n}\Big{(}-\big{(}(y^{\prime}_{m^{\prime}}-f_{W}(x^{\prime}_{m^{\prime}}))\!\cdot\!\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{1}}\big{)}\!+\!\big{(}(\tilde{y}^{\prime}_{m^{\prime}}-f_{W}(\tilde{x}^{\prime}_{m^{\prime}}))\!\cdot\!\mathbbm{1}_{f_{W}(\tilde{x}^{\prime}_{m^{\prime}})\in I_{1}}\big{)}\Big{)}+$
[…]
$+\frac{t(-1)^{v_{B}}}{n}\cdot\Big{(}-\big{(}(y^{\prime}_{m^{\prime}}-f_{W}(x^{\prime}_{m^{\prime}}))\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{B}}\big{)}+\big{(}(\tilde{y}_{m^{\prime}}-f_{W}(\tilde{x}_{m^{\prime}}))\cdot\mathbbm{1}_{f_{W}(\tilde{x}_{m^{\prime}})\in I_{B}}\big{)}\Big{)}$ (33)
$\leq\frac{2t}{n},$ (34)
The last inequality is derived as follows: By definition of the binning, each data point is assigned to exactly one bin. Consequently, for the input $x^{\prime}_{m^{\prime}}$, one indicator from the set $\{\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{i}}\}_{i=1}^{B}$ is non-zero, we denoted the corresponding index $b$. Thus, $\mathbbm{1}_{f_{w}(x^{\prime}_{m^{\prime}})\in I_{b}}\neq 0$, and $\mathbbm{1}_{f_{w}(x^{\prime}_{m^{\prime}})\in I_{b^{\prime}\neq b}}=0$. A similar discussion applies to the input $\tilde{x}_{m^{\prime}}$ with the non-zero index denoted as $\tilde{b}$, indicating $\mathbbm{1}_{f_{W}(\tilde{x}^{\prime}_{m^{\prime}})\in I_{\tilde{b}}}\neq 0$ and $\mathbbm{1}_{f_{w}(x^{\prime}_{m^{\prime}})\in I_{b^{\prime}\neq\tilde{b}}}=0$. Note that $b$ and $\tilde{b}$ can be either identical or different. Therefore, although Eq. (33) contains $2B$ indicator functions, at most only two are non-zero.
Combined with the fact that $|y^{\prime}_{m^{\prime}}-f_{w}(x^{\prime}_{m^{\prime}})|\leq 1$, we obtain Eq. (34). Note that by Assumption 3, $\{f_{w}(x_{m})\}_{m=1}^{n}$ in $x_{m}\in S^{\prime}$ takes the distinct values almost surely and in the above discussion, we do not consider the case when $b/B=f_{w}(x_{m})$ for some $b$ holds, which means that the predicted probability is just the value of the boundary of bins.
Combined with Lemma 1, we have that
$\mathbb{E}_{S^{\prime}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}$
$\leq\sum_{v_{1},\dots,v_{B}=0,1}e^{(t^{2}/8)n\left(\frac{2}{n}\right)^{2}}$
$=2^{B}e^{\frac{t^{2}}{2n}}.$ (35)
When we do not assume that Assumption 3, there may be a possibility that $b/B=f_{w}(x_{m})$ for some $b$ holds, which means that the predicted probability is just the value of the boundary of bins. Then in Eq. (33), at most only four indicator functions are not zero. This results in a worse bound
$\sup_{\{z_{m}\}_{m=1},\tilde{z}_{m}\in\mathcal{Z}}\sum_{i=1}^{B}t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(z_{m}^{\prime})\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})-\frac{1}{n}\sum_{m\neq m^{\prime}}^{n}l_{i}(z_{m}^{\prime})-\frac{1}{n}l_{i}(\tilde{z}_{m^{\prime}})\right]$
$\leq\frac{4t}{n},$ (36)
and results in
$\mathbb{E}_{S^{\prime}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}\leq 2^{B}e^{\frac{2t^{2}}{n}}.$ (37)
We define the upper bound of the exponential moment as
[…]
From Eq. (30), for any $w$ we have
$\mathbb{E}_{S}e^{t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|-f(t,n)}\leq 1.$ (39)
Then we take the expectation with respect to the prior
$\mathbb{E}_{\pi}\mathbb{E}_{S}e^{t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|-f(t,n)}\leq 1.$
Using the Fubini theorem, we change the order of expectations,
$\mathbb{E}_{S}\mathbb{E}_{\pi}\left[{\rm e}^{t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|-f(t,n)}\right]\leq 1.$
and use the Donsker–Varadhan lemma, we have
$\mathbb{E}_{S}\left[{\rm e}^{\sup_{\rho\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\rho}t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|-\mathrm{KL}(\rho\|\pi)-f(t,n)}\right]\leq 1.$
where $\mathcal{P}(\mathcal{W})$ be the set of all probability distributions on $\mathcal{W}$.
By using the Markov inequality with Chernoff-bounding technique, we have
$P\left(\sup_{\rho\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\rho}t|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|-\mathrm{KL}(\rho\|\pi)-f(t,n)>\log\frac{1}{\varepsilon}\right)\leq\varepsilon.$
By rearranging the inequality
$P\left(\exists\rho\in\mathcal{P}(\mathcal{W})\text{, }\mathbb{E}_{w\sim\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|>\frac{\mathrm{KL}(\rho\|\pi)+f(t,n)+\log\frac{1}{\varepsilon}}{t}\right)\leq\varepsilon.$ (40)
By substituting Eq. (38) when without Assumption 3, we have
$P\left(\exists\rho\in\mathcal{P}(\mathcal{W})\text{, }\mathbb{E}_{w\sim\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|>\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\frac{2t^{2}}{n}+\log\frac{1}{\varepsilon}}{t}\right)\leq\varepsilon.$ (41)
As for the ECE of the multiclass setting, from Eq. (24) and Eq. (26), we can proceed the proof exactly the same way as the binary setting, which results in the same upper bound. ∎
A.2.3 Proof of Theorem 2 (Generalization of the ECE under the training dataset)
Now we present the proof of Theorem 2
Theorem 6.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. For both binary and multiclass settings, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{W},S_{\mathrm{te}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (42)
When Assumption 3 holds, we have
$\mathbb{E}_{\rho}|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{W},S_{\mathrm{te}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{n}}{\lambda}.$ (43)
Proof.
To derive the PAC Bayes bound, we transform the loss
$|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{W},S_{\mathrm{te}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|$
$=\left|\mathbb{E}_{S_{\mathrm{te}}}\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{te}}}\left[(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{W}(X)\in I_{i}}\right]\right|-\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{tr}}}(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{W}(X)\in I_{i}}\right|\right|$
$\leq\mathbb{E}_{S_{\mathrm{te}}}\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right|,$ (44)
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$.
We then evaluate the exponential moment as
$\ \mathbb{E}_{S_{\mathrm{tr}},\pi}e^{t|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{W},S_{\mathrm{te}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|}$ $\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}e^{t\mathbb{E}_{S_{\mathrm{te}}}\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right|}$
$\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right|}.$ (45)
By setting $g(i,S_{\mathrm{tr}},S_{\mathrm{te}})\coloneqq\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)$, we have
$\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}|g(i,S_{\mathrm{tr}},S_{\mathrm{te}})|}$ $=\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}\prod_{i=1}^{B}e^{t|g(i,S_{\mathrm{tr}},S_{\mathrm{te}})|}$
$\leq\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}\prod_{i=1}^{B}\left(e^{tg(i,S_{\mathrm{tr}},S_{\mathrm{te}})}+e^{-tg(i,S_{\mathrm{tr}},S_{\mathrm{te}})}\right)$
$\leq\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}\sum_{v_{1},\dots,v_{B}=0,1}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}g(i,S_{\mathrm{tr}},S_{\mathrm{te}})}$ (46)
$=\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}g(i,S_{\mathrm{tr}},S_{\mathrm{te}})}$
$=\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right]},$ (47)
where $\sum_{v_{1},\dots,v_{B}=0,1}$ is all the combinations that will be generated by expanding $\prod_{i=1}^{B}$ in Eq. (46) and it has $2^{B}$ combinations.
We would like to upper bound $\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right]}$ using Lemma 1. For that purpose, here we evaluate $c_{i}$s of Lemma 1. We can upper bound it completely in the same way as Eq. (34). For that purpose, we define $S=S_{\mathrm{te}}\cup S_{\mathrm{tr}}$, $\tilde{g}(S)\coloneqq t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\frac{1}{n}\sum_{Z\in S_{\mathrm{te}}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right]$. Note that $S$ consists of $2n$ random variables. We remark that $\mathbb{E}_{S}[\tilde{g}(S)]=0$ conditioned on $w$, thus, it satisfies the condition of Lemma 1. Then We define $S^{\prime}$ in which we replace single $z_{m}\in S$ with $\tilde{z}_{m}\in\mathcal{Z}$. Then with Assumption 3, we have
$\sup_{S,\tilde{z}_{m}\in\mathcal{Z}}\tilde{g}(S)-\tilde{g}(S^{\prime})\leq\frac{2t}{n},$ (48)
which is followed by a discussion in Eq. (34). Differently from Eq. (34), there exist $4B$ bins, but since we only replace single data point $z_{m}$, as discussed in Eq. (34), only two indicator function is not zero. This results in the above upper bound.
Combined with Lemma 1, we have that
$\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{\tilde{g}(S)}\leq\sum_{v_{1},\dots,v_{B}=0,1}e^{(t^{2}/8)(2n)\left(\frac{2}{n}\right)^{2}}=2^{B}e^{\frac{t^{2}}{n}}.$ (49)
We can derive the upper bound without Assumption 3 in the same way as Eqs. (38) and (48),
$\sup_{S,\tilde{z}_{m}\in\mathcal{Z}}\tilde{g}(S)-\tilde{g}(S^{\prime})\leq\frac{4t}{n},$ (50)
Combined with Lemma 1, we have that
$\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\mathbb{E}_{S_{\mathrm{tr}},S_{\mathrm{te}}}e^{\tilde{g}(S)}\leq\sum_{v_{1},\dots,v_{B}=0,1}e^{(t^{2}/8)(2n)\left(\frac{4}{n}\right)^{2}}=2^{B}e^{\frac{4t^{2}}{n}}.$ (51)
In conclusion, we have
[…]
Then we can proceed the proof of PAC Bayesian bound in the same way as Eq. (39) and repeat the same derivation. This concludes the proof.
As for the ECE of the multiclass setting, from Eq. (24) and Eq. (26), we can proceed the proof exactly the same way as the binary setting, which results in the same upper bound. ∎
A.3 Proof of Theorem 3 (Estimation bias under the training dataset)
First, we present the complete statement again;
Theorem 7.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda},$ (53)
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (54)
Proof.
Binary setting
We start from the binary classification setting. As stated in the proof outline, we decompose the estimation bias into two biases as follows:
$|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!\leq|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|+|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|,$ (55)
where
$f_{\mathcal{I}}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}.$ (56)
In the above, the second term, $\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ has already been proved in Theorem 5. So we only need to bound the first term. As for the first term, we show that
$|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)]-\mathbb{E}[Y|f_{\mathcal{I}}(X)]|+\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|.$ (57)
holds. First, we can show that
$\mathrm{CE}(f_{\mathcal{I}})\leq\mathrm{CE}(f_{w}),$ (58)
by Jensen inequality with respect to the conditional expectation. This has been proved in Proposition 3.3 in Kumar et al. [18]. Next, we upper bound $\mathrm{CE}(f_{w})$ as follows
$\mathrm{CE}(f_{w})$ $=\mathbb{E}\left[|\mathbb{E}[Y|f_{w}(X)]-f_{w}(X)|\right]$
$=\sum_{i=1}^{B}\mathbb{E}[\mathbbm{1}_{f_{w}(X)\in I_{i}}\cdot|\mathbb{E}[Y|f_{w}(X)]-f_{w}(X)|]$
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f_{w}(X)]-f_{w}(X)||f_{w}(X)\in I_{i}]$
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f_{w}(X)]-\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]$
$\quad+\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]-f_{w}(X)||f_{w}(X)\in I_{i}]$
$\leq\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}||\mathbb{E}[Y|f_{w}(X)]-\mathbb{E}[Y|f_{w}(X)\in I_{i}]|$
$\quad+\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}|\mathbb{E}[Y|f_{w}(X)\in I_{i}]-\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]|$
$\quad+\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}[|\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]-f_{w}(X)||f_{w}(X)\in I_{i}],$
where the second term is $\mathrm{CE}(f_{\mathcal{I}})$, we finished the proof of Eq. (57).
Then we can show that
$|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)]-\mathbb{E}[Y|f_{\mathcal{I}}(X)]|+\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|\leq\frac{L}{B}+\frac{1}{B},$ (60)
where we used the Lipschitz continuity of the function and the fact that we split the function with equal width $1/B$. Then using Eq. (55) and above result, we have
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ $\leq\mathbb{E}_{\rho}|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|+\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$
$\leq\frac{L}{B}+\frac{1}{B}+\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|,$ (61)
and use Theorem 5 for the second term, this concludes the proof for the binary setting. ∎
Proof.
Multiclass setting
Next, for the multiclass setting, similarly to the binary case, we introduce the conditional expectation defined as;
$f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)_{C}\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}.$ (62)
As for the multiclass setting, similarly to the binary case, we have
$|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!\leq|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}})|+|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|.$ (63)
Then similarly to Eq. (57) in the binary classification, we can show that
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)_{C}]-\mathbb{E}[Y|f_{\mathcal{I}}^{C}(X)]|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|$ (64)
by the almost identical proof. We then upper bound them
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)_{C}]-\mathbb{E}[Y|f_{\mathcal{I}}^{C}(X)]|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|\leq\frac{L}{B}+\frac{1}{B}$ (65)
using the Lipschitz continuity and the property of the binning. We substitute this into Eq. (63), we have
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!$
$\leq\mathbb{E}_{\rho}|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}})|+\mathbb{E}_{\rho}|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$
$\leq\frac{L}{B}+\frac{1}{B}+\mathbb{E}_{\rho}|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ (66)
Then using Theorem 5 for the second term, this concludes the proof. ∎
A.4 Limitation under the Hölder continuity
Here we discuss the assumption of Lipschitz continuity. It is known that by assuming the higher order smoothness, such as the $\beta$-Hölder continuity, the bias of the nonparametric estimation decreases [30], particularly, the lower bound is $\mathcal{O}(n^{-\frac{\beta}{2\beta+1}})$.
However, in the binning scheme, we cannot improve the optimal order from $\mathcal{O}(n^{-1/3})$ due to the binning bias. This is because that in Eq. (60), which is appearing the proof of the estimation bias, there is the error term $\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|$. This term is upper bounded by $1/B$ since we consider that the bins are allocated with equal width. Thus, we cannot improve the estimation bias order owing to this error term, which remains $1/B$ even under the Hölder continuity assumption. Thus, the binning method cannot utilize the smoothness of the underlying data distribution.
A.5 Proof of Theorem 4 (Estimation bias for the multiclass setting)
The upper bound can be derived similarly to the binary classification setting in Appendix A.3; 1) we first transform the $ECE_{K}$ and $CE_{K}$, 2) decompose the estimation bias to the generalization and binning bias, 3) bound those two terms using the generalization error analysis and the binning definitions.
[…]
ECE transformation
We first derive the ECE transformation; for the $ECE_{K}$ under the training dataset, we have
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})=\sum_{i=1}^{B}\frac{\#B_{i}}{n}\|\bar{f}_{i,S_{\mathrm{tr}}}-\bar{y}_{i,S_{\mathrm{tr}}}\|=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|.$ (67)
Next, we define the conditional expectation as
$f_{\mathcal{I}}(x)\coloneqq\sum_{i=1}^{B}f_{I_{i}}(x)\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}=\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}.$ (68)
Then by repeating the exactly same proof for Eq. (21), the following relation holds
$\mathrm{CE}_{K}(f_{\mathcal{I}})=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|.$ (69)
Again, we consider the following decomposition for the estimation bias
$|\mathrm{CE}_{K}(f_{w})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|\!\leq|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|+|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|,$ (70)
We then separately upper bound the above two terms.
Generalization of ECE
First we focus on the second term, $|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|$, and we upper bound it by the PAC Bayes bound;
Theorem 8.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$ and $f_{w}(X)$ has the probability density over $[0,1]^{K}$. Then for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|CE_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{K^{2}\lambda^{2}}{2n}}{\lambda}.$ (71)
Proof.
The proof strategy is almost identical to Appendix A.2.2.
First, we transform the generalization gap by using Eqs. (67) and  (69). First note that
$\mathrm{CE}_{K}(f_{\mathcal{I}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|$ (72)
and
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|$ (73)
Thus, we have
$|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|$
$=\left|\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|-\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|\right|$
$\leq\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|,$ (74)
where we used the triangle inequality $|\|a\|-\|b\||\leq\|a-b\|$ for the first inequality and set $l_{i}(z)=(\hat{y}-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$. We then evaluate the exponential moment as follows;
$\mathbb{E}_{S_{\mathrm{tr}},\pi}e^{t|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|}$ $\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|}$
$\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)_{k}-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)_{k}|}$ (75)
where $l_{i}(Z)_{k}$ is the $k$-th dimension of $l_{i}(z)$.
By setting $g(i,k,S_{\mathrm{tr}})\coloneqq\mathbb{E}_{\tilde{Z}}l_{i}(\tilde{Z})_{k}-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)_{k}$, we have
$\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|g(i,k,S_{\mathrm{tr}})|}$ $=\mathbb{E}_{S_{\mathrm{tr}}}\prod_{i=1}^{B}\prod_{k=1}^{K}e^{t|g(i,k,S_{\mathrm{tr}})|}$
$\leq\mathbb{E}_{S_{\mathrm{tr}}}\prod_{i=1}^{B}\prod_{k=1}^{K}\left(e^{tg(i,k,S_{\mathrm{tr}})}+e^{-tg(i,k,S_{\mathrm{tr}})}\right)$
$\leq\mathbb{E}_{S_{\mathrm{tr}}}\sum_{v_{1},\dots,v_{BK}=0,1}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}(-1)^{v_{i}}g(i,k,S_{\mathrm{tr}})}$ (76)
$=\sum_{v_{1},\dots,v_{BK}=0,1}\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}(-1)^{v_{i}}g(i,k,S_{\mathrm{tr}})}$
$=\sum_{v_{1},\dots,v_{BK}=0,1}\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}(-1)^{v_{i}}(\mathbb{E}_{\tilde{Z}}l_{i}(\tilde{Z})_{k}-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)_{k})}$ (77)
where $\sum_{v_{1},\dots,v_{BK}=0,1}$ is all the combinations that will be generated by expanding $\prod_{i=1}^{BK}$ in Eq. (46) and it has $2^{BK}$ combinations.
We would like to upper bound the exponential moment using Lemma 1. For that purpose, here we evaluate $c_{i}$s of Lemma 1. We can upper-bound it completely in the same way as the binary case. Then with Assumption that $f_{w}(x)$ is absolutely continuous, $f_{w}(x)$ takes the distinct values almost surely. Thus, we have
$\sup_{\{z_{m}\}_{m=1},\tilde{z}_{m}\in\mathcal{Z}}\sum_{i=1}^{B}\sum_{k=1}^{K}t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})_{k}-\frac{1}{n}\sum_{m=1}^{n}l_{i}(z_{m}^{\prime})_{k}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad-t(-1)^{v_{i}}\cdot\left[\mathbb{E}_{Z^{\prime}}l_{i}(Z^{\prime})_{k}-\frac{1}{n}\sum_{m\neq m^{\prime}}^{n}l_{i}(z_{m}^{\prime})_{k}-\frac{1}{n}l_{i}(\tilde{z}_{m^{\prime}})_{k}\right]$
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\sum_{i=1}^{B}\sum_{k=1}^{K}\frac{t(-1)^{v_{i}}}{n}\cdot\left[-l_{i}(z^{\prime}_{m^{\prime}})_{k}+l_{i}(\tilde{z}_{m^{\prime}})_{k}\right]$
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\sum_{k=1}^{K}\frac{t(-1)^{v_{1}}}{n}\cdot\Big{(}-\big{(}(\hat{y}^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{1}}\big{)}+\big{(}(\hat{\tilde{y}}^{\prime}_{m^{\prime},k}-f_{W}(\tilde{x}^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}^{\prime}_{m^{\prime}})\in I_{1}}\big{)}\Big{)}+$
[…]
$+\frac{t(-1)^{v_{B}}}{n}\cdot\Big{(}-\big{(}(\hat{y}^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{B}}\big{)}+\big{(}(\hat{\tilde{y}}_{m^{\prime},k}-f_{W}(\tilde{x}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}_{m^{\prime}})\in I_{B}}\big{)}\Big{)}$ (78)
$\leq\sum_{k=1}^{K}\frac{2t}{n}\leq\frac{2Kt}{n},$ (79)
which is followed by a discussion in Eq. (34). Combined with Lemma 1, we have that
$\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|g(i,S_{\mathrm{tr}})|}$ $\leq\sum_{v_{1},\dots,v_{BK}=0,1}e^{(t^{2}/8)(n)\left(\frac{2K}{n}\right)^{2}}=2^{BK}e^{\frac{K^{2}t^{2}}{2n}}.$ (80)
Thus, the exponential moment is upper-bounded and we define
$f(t,n)\coloneqq 2^{BK}e^{\frac{K^{2}t^{2}}{2n}}$ (81)
We then repeat the Chernoff-bounding technique of the proof of the PAC-Bayes bound derivation in Appendix A.2.2, and substitute $f(t,n)$ into Eq. (40), we have the following PAC-Bayes bound; for any prior $\pi$ and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}|CE_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{K^{2}\lambda^{2}}{2n}}{\lambda}.$ (82)
Bias analysis of the binning method
Next, we upper bound $|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|$, which is the first term of Eq. (70). By repeating exactly the same procedure in Eq. (57), we can show that
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|\leq\mathbb{E}\|\mathbb{E}[\hat{Y}|f_{w}(X)]-\mathbb{E}[\hat{Y}|f_{\mathcal{I}}(X)]\|+\mathbb{E}\|f_{w}(X)-f_{\mathcal{I}}(X)\|.$ (83)
We can bound them using the definition of the binning construction; recall that in the current binning construction, we split each dimension of $[0,1]^{K}$ with $B^{\prime}$ bins with equal width. Thus $|f_{w}(X)_{k}-f_{\mathcal{I}}(X)_{k}|\leq 1/B^{\prime}$ for all $k\in[K]$. This implies that
$\mathbb{E}\|f_{w}(X)-f_{\mathcal{I}}(X)\|\leq\mathbb{E}(\sum_{k=1}^{K}|f_{w}(X)_{k}-f_{\mathcal{I}}(X)_{k}|)\leq\sum_{k=1}^{K}\frac{1}{B^{\prime}}=\frac{K}{B^{\prime}}$ (84)
Then we use the fact that $B=(B^{\prime})^{L}$, we have that
$\mathbb{E}\|f_{w}(X)-f_{\mathcal{I}}(X)\|\leq\frac{K}{B^{\frac{1}{K}}}$ (85)
Using the Lipschitz continuity for the first, we have
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|$ $\leq\mathbb{E}||\mathbb{E}[\hat{Y}|f_{w}(X)]-\mathbb{E}[\hat{Y}|f_{\mathcal{I}}(X)]|+\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|$
$\leq\frac{KL}{B^{\prime}}+\frac{K}{B^{\prime}}=\frac{K(1+L)}{B^{\frac{1}{K}}}.$ (86)
Conclusion
Finally, taking the expectation by $\rho$ in Eq. (70), we upper bound the first and second term using Eq. (A.5) and Eq. (82), we obtain the theorem.
A.6 Discussion about the assumption of Theorem 4
In Theorem 4, we assumed that $f_{w}(x)$ has the probability density in $[0,1]^{K}$. As discussed in Appendix A.2.2, in the binary classification and TCE analysis, this assumption is not required, and if we assume this, our bound results in the improvement of the coefficient, not affect the order of the bound.
On the other hand, when considering the generalization of $ECE_{K}$, this assumption is inevitable to circumvent the curse of dimensionality especially in the generalization error analysis. Here we discuss how Theorem 4 changes when eliminating this assumption. If we do not assume this assumption, in the proof of Theorem 4, the exponential moment is replaced to
$\mathbb{E}_{S_{\mathrm{tr}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|g(i,S_{\mathrm{tr}})|}$ $\leq\sum_{v_{1},\dots,v_{BK}=0,1}e^{(t^{2}/8)(n)\left(\frac{2K(2^{K})}{n}\right)^{2}}=2^{BK}e^{\frac{4^{K}K^{2}t^{2}}{2n}},$ (87)
which is the surprisingly worse upper bound compared with the upper bound of the exponential moment shown in Eq. (80) with the assumption. Then the obtained PAC Bayesian bound becomes
$\mathbb{E}_{\rho}|CE_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+BK\log 2+\log\frac{1}{\varepsilon}+\frac{4^{K}K^{2}\lambda^{2}}{2n}}{\lambda}.$ (88)
which has the new coefficient $4^{K}$ and it is significantly larger than that of Theorem 4.
This is caused by the evaluation of the coefficient of the Lemma 1. In the derivation of Eq. (A.5), if we do not assume that $f_{w}$ has the probability density, then there is a possibility that the training data is allocated to the point at the exact grid points of the hypercube constructed by the bins. The grid point is the intersection of $2^{K}$ small regions, and thus, the $c_{i}$s of Lemma 1 becomes exponentially large. Owing to this, the coefficient of Eq. (A.5) is significantly larger compared with the setting when we assume that $f_{w}$ has the probability density, where we do not need to consider such a worst case. In this way, the order of the bias significantly improves by assuming that $f_{w}$ has the probability density.
Appendix B Proofs and discussion of Section 4
We remark that recalibration takes $\eta_{V}:\Delta^{K}\to\Delta^{K}$, and this is clear for the multiclass setting since $f_{w}:\mathcal{X}\to\Delta^{K}$. In the case of the binary setting, $f_{w}:\mathcal{X}\to[0,1]$. So, we consider that the input to $\eta_{v}$ is $(f_{w}(x),1-f_{w}(x))\in\Delta^{2}$. Then we can treat the recalibration in a unified way.
B.1 Proofs for Corollary 1 and 2
When we recalibrate the model, we fix $w$. Thus under the $f_{w}$, the given recalibration data $S_{\mathrm{re}}=\{(X_{m},Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$, it is transformed into $\tilde{S}_{\mathrm{re}}=\{(f_{w}(X_{m}),Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$ and samples in $\tilde{S}_{\mathrm{re}}$ are i.i.d by definition. Then we apply our developed PAC-Bayes bounds in Appendix A for this new dataset.
As for Corollary 1, we can prove it exactly in the same way as shown in Appendix A.2.3. Simply replacing $S_{\mathrm{tr}}$ with $\tilde{S}_{\mathrm{re}}$ in Appendix A.2.3 and apply it to the standard derivation of PAC-Byaeisan bound shown in Eq. (39). This concludes the proof.
As for Corollary 2, we first provide the complete statement;
Corollary 3.
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$. Under Assumption 2, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda},$ (89)
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (90)
We can prove this exactly in the same way as shown in Appendix A.3. Simply replacing $S_{\mathrm{tr}}$ with $\tilde{S}_{\mathrm{re}}$ in Appendix A.3. This concludes the proof.
B.2 Discussion about the conditional KL divergence
In Corollary 1 and 2, we fixed $w$. If we take the expectation with respect to $W$, then the bound of Corollary 1 become as follows;
Corollary 4.
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$. Under Assumption 2, conditioned on $S_{\mathrm{tr}}=s_{\mathrm{tr}}$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\rho(W|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{V}(f_{W}),S_{\mathrm{re}},\mathrm{ECE})\leq\frac{\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}[\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})]+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (91)
where $\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}[\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})]$ represents the conditional KL divergence.
Proof.
The proof is almost the same as Appendix A.2.3. The difference is how we utilize the Chernoff-bounding technique. Here we only show the different parts. After obtaining the upper bound of the exponential moment $f(t,n_{\mathrm{re}})$ as in Eq. (52), we substitute this into Eq. (39), which can be written as follows; for any $w$ and $v$ we have
$\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$ (92)
Then we take the expectation with respect to the prior $\tilde{\pi}$ and $\rho(w|s_{\mathrm{tr}})$,
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$
Using the Fubini theorem, we change the order of expectations,
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1,$
and $\rho(w|s_{\mathrm{tr}})$ and $S_{\mathrm{re}}$ are independent thus
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1.$
Then, use the Donsker–Varadhan lemma, we have
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$
By taking the Jensen inequality, we have
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (93)
By taking the swap about $\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}$ and $\sup$, we have
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (94)
After this, we simply follow the derivation of the standard PAC-Bayes derivation in Appendix A.3. ∎
B.3 Discussion for Section 4.2
First, the upper bound of the ECE is obtained by
$\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|\hat{Y}_{C}-\eta_{v}(f_{w}(X))_{C}|^{2}\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\hat{Y}-\eta_{v}(f_{w}(X))\|_{2}^{2}$ (95)
where $\hat{Y}_{C}$ is the $C$-th dimension of the one-hot encoding of the label $Y$ and the first inequality is followed by Hölder inequality and the Jensen inequality. As for the binary classification, by Hölder inequality and the Jensen inequality, we have
$\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|Y-\eta_{v}(f_{w}(X))_{1}|^{2},$ (96)
since the recalibrated function is $\eta_{v}(f_{w}(X))\in\Delta^{2}$ and its 1st dimension represents the predicted probability of the label is $1$.
Then from the estimation bias bound in Eq. (14), we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{TCE}(\eta_{v}(f_{w}))$
$\leq\mathbb{E}_{\tilde{\rho}}\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})+\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}$
$\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2}+\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda},$ (97)
holds for any $\lambda>0$. Thus, the objective function of Eq. (15) is the upper bound of the TCE in the PAC-Bayes bound. Clearly, a similar statement holds for $CE$ in the binary setting.
Next, we discuss the objective function of Eq. (16). This objective function can guarantee the performance of test accuracy and TCE simultaneously as follows; Assume that $l_{\mathrm{acc}}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and define the training loss as $\hat{L}_{\mathrm{acc}}(V,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}^{n_{\mathrm{tr}}}l_{\mathrm{acc}}(Y,\eta_{v}(f_{w}(X)))$ and the (expected) test loss as $L_{\mathrm{acc}}(v,\mathcal{D})\coloneqq\mathbb{E}_{Z}l_{\mathrm{acc}}(Y,\eta_{v}(f_{w}(X)))$. Under this setting, we have
Corollary 5.
Assume that $n_{\mathrm{te}}=n_{\mathrm{re}}=n$ and . Under Assumption 2, conditioned on $W=w$, for any fixed prior $\tilde{\pi}$, which is independent of $S_{\mathrm{re}}$, and posterior distribution $\tilde{\rho}$ over $V$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{re}}$, we have
$\mathbb{E}_{\tilde{\rho}}(L_{\mathrm{acc}}(V,\mathcal{D})+\mathrm{TCE}(\eta_{V}(f_{w}))$ $\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}(l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2})$
$\quad+\frac{1+L}{B}+\frac{2\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\frac{33\lambda^{2}}{8n_{\mathrm{re}}}+2\log\frac{2}{\varepsilon}}{\lambda},$ (98)
Proof.
From Eq. (1), which is the standard PAC-Bayes bound for $[0,1]$ bounded loss function, we have
$\mathbb{E}_{\tilde{\rho}}L_{\mathrm{acc}}(V,\mathcal{D})\leq\mathbb{E}_{\tilde{\rho}}\hat{L}_{\mathrm{acc}}(V,S_{\mathrm{tr}})+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{8n_{\mathrm{re}}}}{\lambda}.$ (99)
By considering the union bound with Eq. (99) and Eq. (B.3), setting $\varepsilon$ as $\varepsilon/2$ in each high probability bound, we obtain the result. ∎
Clearly, a similar statement holds for $CE$ in the binary setting.
Thus, the objective function Eq. (16) optimizes the upper bound of Corollary 5, and thus, it naturally guarantees the test accuracy and TCE simultaneously.
Table 2: Datasets used in our experiments
Dataset Classes Train data ($n_{\mathrm{tr}}$) Recalibration data ($n_{\mathrm{re}}$) Test data ($n_{\mathrm{te}}$)
KITTI [8] $2$ $16000$ $1000$ $8000$
PCam [31] $2$ $22768$ $1000$ $9000$
MNIST [19] $10$ $60000$ $1000$ $9000$
CIFAR-100 [15] $100$ $50000$ $1000$ $9000$
Appendix C Details of experimental settings
In this section, we summarize the detail information of our numerical experiments in Section 6. Our CIFAR-100 experiments were conducted on NVIDIA GPUs with $32$GB memory (NVIDIA DGX-1 with Tesla V100 and DGX-2). For the other experiments, we used CPU (Apple M1) with $16$GB memory.
Datasets and models:
We show the details of the datasets and the numbers of training, recalibration, and test data in Table 2. For the models, we used XGBoost [5], Random Forests [3], and a $1$-layer neural network (NN) for the KITTI, PCam, and MNIST experiments. For the CIFAR-100 experiments, we used AlexNet [16], WideResNet [37], DenseNet-BC-190 [13], and ResNeXt-29 [34]. XGBoost, Random Forests, and the $1$-layer NN were trained by adapting the code from Wenger et al. [32]  [^3] https://github.com/JonathanWenger/pycalib. Models used in the CIFAR-100 experiments were obtained from the GitHub project page: pytorch-classification (https://github.com/bearpaw/pytorch-classification).
Recalibration algorithms:
In the multiclass classification experiments, the standard method of temperature scaling [10] and the recently developed GP calibration [32] that has shown good performance were used as baselines. As a baseline method for binary classification, we additionally adopted Platt scaling [24], isotonic regression [36], Beta calibration [17], and Bayesian binning into quantiles (BBQ) [23]. We also adopted the GP calibration with mean approximation [32] in the additional experiments in Appendix D.
For our PBR method, we set the Gaussian distribution as both the posterior and prior, as in Wenger et al. [32]. We also followed the settings for training strategy, hyperparameters, and optimizers in Wenger et al. [32] to ensure a fair comparison. PBR has an optimizable parameter $1/\lambda$ in addition to the posterior parameter; thus, we selected it using grid search from the following candidates: $0\}$ , 0.01 , 0.1 , 0.25 , 0.5 , 0.75 , 0.9 , 1.0 }. The mean approximation strategy in Wenger et al. [32] can be applied for PBR. Thus, we reported the results of PBR with mean approximation (denoted as PBR app.) in Appendix D.
Performance evaluation:
We measured the predictive accuracy and the ECE estimated by binning, with $B=\lfloor n_{\mathrm{re}}^{1/3}\rfloor$, following our theoretical findings in Corollary 2. We conducted $10$-fold cross-validation for recalibration function training and reported the mean and standard deviation of these two performance metrics. We also used $J=100$ Monte Carlo samples from posterior $\tilde{\rho}$ to obtain $V$.
Appendix D Additional experimental results
Here are additional experimental results that could not be included in the main part of this paper.
We first show all the experimental results regarding the correlation between the KL-divergence and our bound in Figures 2 and 3. We observed that the KL-divergence and generalization gap are well correlated in terms of Pearson’s and Kendall’s rank-correlation coefficients, especially in the multiclass classification tasks (MNIST and CIFAR-100). However, for the binary classification experiments, no strong correlation was observed. We conjecture that this is caused by the estimation bias of ECE through binning, which has a very slow convergence rate of $\mathcal{O}(n_{\mathrm{re}}^{1/3})$ under our optimal $B$. Since we use $n_{\mathrm{re}}=1000$, the noise due to the estimation bias can be $\mathcal{O}(1/10)$, which is larger than the ECE gap values in the binary classification experiments. Therefore, it seems that the correlation is buried in the noise.
We also show the results of comparisons with all baseline methods in Tables 3-6. These results confirm that our method almost consistently improves GP-based recalibration methods.
Figure 2: KL vs ECE gap on KITTI (binary), PCam (binary), and MNIST (Multiclass) dataset per various regularize parameters ($\alpha$).
Figure 3: KL vs ECE gap on CIFAR-100 (Multiclass) dataset per various regularize parameters ($\alpha$). For the AlexNet experiments, the x-axis range has been limited in the figure to better illustrate the data due to the observation of very large KL values for the case where $\alpha=0$.
Table 3: KITTI (binary)
Model Method ECE (mean $\pm$ std) Accuracy (mean $\pm$ std)
XGBoost Uncalibration $0.0119\pm 0.0004$ $0.9599\pm 0.0005$
GP calibration [32] $0.0060\pm 0.0015$ $0.9657\pm 0.0007$
GP calibration (app.) [32] $0.0067\pm 0.0017$ $0.9657\pm 0.0008$
PBR (ours) $0.0070\pm 0.0019$ $0.9657\pm 0.0006$
PBR total (ours) $\mathbf{0.0045\pm 0.0016}$ $0.9659\pm 0.0005$
PBR app. (ours) $0.0072\pm 0.0016$ $0.9655\pm 0.0006$
PBR app. total (ours) $0.0046\pm 0.0016$ $0.9659\pm 0.0005$
Platt scaling [24] $0.0076\pm 0.0018$ $0.9624\pm 0.0006$
Isotonic regression [36] $0.0079\pm 0.0023$ $0.9646\pm 0.0011$
Beta calibration [17] $0.0047\pm 0.0013$ $0.9658\pm 0.0005$
BBQ [23] $0.0071\pm 0.0025$ $0.9602\pm 0.0061$
Temperature scaling [10] $0.0090\pm 0.0018$ $0.9599\pm 0.0005$
random forest Uncalibration $0.0734\pm 0.0009$ $0.9638\pm 0.0008$
GP calibration [32] $0.0086\pm 0.0047$ $0.9628\pm 0.0015$
GP calibration (app.) [32] $0.0072\pm 0.0037$ $0.9628\pm 0.0015$
PBR (ours) $0.0090\pm 0.0048$ $0.9629\pm 0.0017$
PBR total (ours) $0.0068\pm 0.0040$ $0.9629\pm 0.0015$
PBR app. (ours) $0.0084\pm 0.0054$ $0.9629\pm 0.0017$
PBR app. total (ours) $0.0068\pm 0.0039$ $0.9629\pm 0.0016$
Platt scaling [24] $\mathbf{0.0048\pm 0.0030}$ $0.9627\pm 0.0011$
Isotonic regression [36] $0.0082\pm 0.0041$ $0.9610\pm 0.0018$
Beta calibration [17] $0.0067\pm 0.0034$ $0.9629\pm 0.0017$
BBQ [23] $0.1064\pm 0.0502$ $0.8943\pm 0.0018$
Temperature scaling [10] $0.0056\pm 0.0025$ $0.9638\pm 0.0008$
1layer NN Uncalibration $\mathbf{0.0049\pm 0.0004}$ $0.9617\pm 0.0010$
GP calibration [32] $0.0057\pm 0.0030$ $0.9682\pm 0.0006$
GP calibration (app.) [32] $0.0063\pm 0.0029$ $0.9682\pm 0.0006$
PBR (ours) $0.0074\pm 0.0037$ $0.9677\pm 0.0007$
PBR total (ours) $0.0060\pm 0.0032$ $0.9679\pm 0.0007$
PBR app. (ours) $0.0077\pm 0.0036$ $0.9678\pm 0.0007$
PBR app. total (ours) $0.0062\pm 0.0039$ $0.9680\pm 0.0007$
Platt scaling [24] $0.0089\pm 0.0039$ $0.9644\pm 0.0009$
Isotonic regression [36] $0.0081\pm 0.0037$ $0.9671\pm 0.0023$
Beta calibration [17] $0.0061\pm 0.0037$ $0.9679\pm 0.0007$
BBQ [23] $0.0316\pm 0.0732$ $0.9627\pm 0.0038$
Temperature scaling [10] $0.0092\pm 0.0042$ $0.9617\pm 0.0010$
Table 4: PCam (binary)
Model Method ECE (mean $\pm$ std) Accuracy (mean $\pm$ std)
XGBoost Uncalibration $0.0516\pm 0.0009$ $0.8448\pm 0.0008$
GP calibration [32] $0.0135\pm 0.0048$ $0.8467\pm 0.0013$
GP calibration (app.) [32] $0.0140\pm 0.0047$ $0.8467\pm 0.0013$
PBR (ours) $0.0123\pm 0.0067$ $0.8467\pm 0.0012$
PBR total (ours) $0.0122\pm 0.0042$ $0.8463\pm 0.0012$
PBR app. (ours) $0.0124\pm 0.0069$ $0.8470\pm 0.0010$
PBR app. total (ours) $\mathbf{0.0121\pm 0.0049}$ $0.8464\pm 0.0013$
Platt scaling [24] $0.0407\pm 0.0103$ $0.8464\pm 0.0014$
Isotonic regression [36] $0.0167\pm 0.0052$ $0.8446\pm 0.0020$
Beta calibration [17] $0.0130\pm 0.0033$ $0.8470\pm 0.0012$
BBQ [23] $0.0178\pm 0.0079$ $0.8441\pm 0.0063$
Temperature scaling [10] $0.0542\pm 0.0070$ $0.8448\pm 0.0008$
random forest Uncalibration $0.0807\pm 0.0010$ $0.8494\pm 0.0012$
GP calibration [32] $0.0133\pm 0.0029$ $0.8485\pm 0.0017$
GP calibration (app.) [32] $0.0125\pm 0.0028$ $0.8485\pm 0.0016$
PBR (ours) $0.0121\pm 0.0044$ $0.8481\pm 0.0016$
PBR total (ours) $0.0105\pm 0.0030$ $0.8482\pm 0.0017$
PBR app. (ours) $0.0121\pm 0.0045$ $0.8481\pm 0.0018$
PBR app. total (ours) $0.0108\pm 0.0033$ $0.8481\pm 0.0020$
Platt scaling [24] $0.0109\pm 0.0034$ $0.8490\pm 0.0014$
Isotonic regression [36] $0.0159\pm 0.0049$ $0.8470\pm 0.0036$
Beta calibration [17] $0.0105\pm 0.0044$ $0.8486\pm 0.0021$
BBQ [23] $0.0491\pm 0.0092$ $0.8078\pm 0.0083$
Temperature scaling [10] $\mathbf{0.0094\pm 0.0023}$ $0.8494\pm 0.0012$
1layer NN Uncalibration $0.2063\pm 0.0012$ $0.5923\pm 0.0012$
GP calibration [32] $0.0390\pm 0.0055$ $0.6498\pm 0.0026$
GP calibration (app.) [32] $0.0394\pm 0.0048$ $0.6499\pm 0.0026$
PBR (ours) $0.0376\pm 0.0086$ $0.6509\pm 0.0015$
PBR total (ours) $0.0349\pm 0.0088$ $0.6503\pm 0.0022$
PBR app. (ours) $0.0382\pm 0.0087$ $0.6508\pm 0.0014$
PBR app. total (ours) $0.0348\pm 0.0090$ $0.6501\pm 0.0024$
Platt scaling [24] $0.0490\pm 0.0122$ $0.6220\pm 0.0059$
Isotonic regression [36] $0.0224\pm 0.0096$ $0.6476\pm 0.0041$
Beta calibration [17] $0.0274\pm 0.0072$ $0.6481\pm 0.0027$
BBQ [23] $\mathbf{0.0168\pm 0.0079}$ $0.6457\pm 0.0068$
Temperature scaling [10] $0.0256\pm 0.0178$ $0.5923\pm 0.0012$
Table 5: MNIST (Multiclass)
Model Method ECE (mean $\pm$ std) Accuracy (mean $\pm$ std)
XGBoost Uncalibration $0.0036\pm 0.0003$ $0.9785\pm 0.0005$
GP calibration [32] $0.0038\pm 0.0005$ $0.9785\pm 0.0007$
GP calibration (app.) [32] $0.0040\pm 0.0005$ $0.9785\pm 0.0006$
PBR (ours) $\mathbf{0.0035\pm 0.0005}$ $0.9784\pm 0.0007$
PBR total (ours) $0.0037\pm 0.0007$ $0.9785\pm 0.0007$
PBR app. (ours) $0.0038\pm 0.0002$ $0.9784\pm 0.0006$
PBR app. total (ours) $0.0041\pm 0.0007$ $0.9785\pm 0.0007$
Temperature scaling [10] $0.0055\pm 0.0017$ $0.9785\pm 0.0005$
random forest Uncalibration $0.1428\pm 0.0007$ $0.9659\pm 0.0005$
GP calibration [32] $0.0324\pm 0.0036$ $0.9663\pm 0.0011$
GP calibration (app.) [32] $0.0139\pm 0.0032$ $0.9663\pm 0.0009$
PBR (ours) $0.0052\pm 0.0018$ $0.9660\pm 0.0009$
PBR total (ours) $0.0066\pm 0.0029$ $0.9661\pm 0.0012$
PBR app. (ours) $0.0085\pm 0.0041$ $0.9662\pm 0.0009$
PBR app. total (ours) $\mathbf{0.0047\pm 0.0015}$ $0.9662\pm 0.0012$
Temperature scaling [10] $0.0062\pm 0.0009$ $0.9659\pm 0.0005$
1layer NN Uncalibration $0.0164\pm 0.0004$ $0.9760\pm 0.0005$
GP calibration [32] $0.0101\pm 0.0023$ $0.9740\pm 0.0008$
GP calibration (app.) [32] $0.0145\pm 0.0017$ $0.9740\pm 0.0009$
PBR (ours) $0.0137\pm 0.0014$ $0.9759\pm 0.0007$
PBR total (ours) $0.0112\pm 0.0019$ $0.9740\pm 0.0016$
PBR app. (ours) $0.0159\pm 0.0007$ $0.9760\pm 0.0006$
PBR app. total (ours) $0.0148\pm 0.0017$ $0.9742\pm 0.0015$
Temperature scaling [10] $\mathbf{0.0062\pm 0.0026}$ $0.9760\pm 0.0005$
Table 6: CIFAR100 (Multiclass)
Model Method ECE (mean $\pm$ std) Accuracy (mean $\pm$ std)
alexnet Uncalibration $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
GP calibration [32] $0.2548\pm 0.0008$ $0.4373\pm 0.0009$
GP calibration (app.) [32] $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
PBR (ours) $0.2548\pm 0.0007$ $0.4372\pm 0.0009$
PBR total (ours) $0.2548\pm 0.0007$ $0.4373\pm 0.0009$
PBR app. (ours) $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
PBR app. total (ours) $0.2548\pm 0.0007$ $0.4374\pm 0.0008$
Temperature scaling [10] $\mathbf{0.0216\pm 0.0037}$ $0.4374\pm 0.0008$
WRN-28-10-drop Uncalibration $0.0568\pm 0.0009$ $0.8131\pm 0.0011$
GP calibration [32] $0.0567\pm 0.0010$ $0.8129\pm 0.0011$
GP calibration (app.) [32] $0.0569\pm 0.0009$ $0.8130\pm 0.0011$
PBR (ours) $0.0504\pm 0.0054$ $0.8129\pm 0.0011$
PBR total (ours) $\mathbf{0.0349\pm 0.0093}$ $0.8129\pm 0.0011$
PBR app. (ours) $0.0569\pm 0.0007$ $0.8130\pm 0.0012$
PBR app. total (ours) $0.0569\pm 0.0008$ $0.8129\pm 0.0011$
Temperature scaling [10] $0.0374\pm 0.0028$ $0.8131\pm 0.0011$
resnext-8x64d Uncalibration $0.0401\pm 0.0010$ $0.8229\pm 0.0013$
GP calibration [32] $0.0400\pm 0.0010$ $0.8229\pm 0.0014$
GP calibration (app.) [32] $0.0401\pm 0.0009$ $0.8229\pm 0.0013$
PBR (ours) $0.0320\pm 0.0050$ $0.8228\pm 0.0014$
PBR total (ours) $\mathbf{0.0310\pm 0.0082}$ $0.8230\pm 0.0014$
PBR app. (ours) $0.0399\pm 0.0024$ $0.8230\pm 0.0013$
PBR app. total (ours) $0.0402\pm 0.0008$ $0.8230\pm 0.0014$
Temperature scaling [10] $0.0401\pm 0.0019$ $0.8229\pm 0.0013$
resnext-16x64d Uncalibration $0.0405\pm 0.0012$ $0.8231\pm 0.0013$
GP calibration [32] $0.0403\pm 0.0013$ $0.8230\pm 0.0014$
GP calibration (app.) [32] $0.0405\pm 0.0014$ $0.8230\pm 0.0014$
PBR (ours) $\mathbf{0.0305\pm 0.0056}$ $0.8231\pm 0.0013$
PBR total (ours) $0.0321\pm 0.0071$ $0.8230\pm 0.0014$
PBR app. (ours) $0.0384\pm 0.0032$ $0.8230\pm 0.0014$
PBR app. total (ours) $0.0369\pm 0.0033$ $0.8229\pm 0.0014$
Temperature scaling [10] $0.0420\pm 0.0025$ $0.8231\pm 0.0013$
densenet-bc-L190-k40 Uncalibration $0.0639\pm 0.0008$ $\mathbf{0.8230\pm 0.0007}$
GP calibration [32] $0.0639\pm 0.0008$ $0.8229\pm 0.0007$
GP calibration (app.) [32] $0.0640\pm 0.0008$ $0.8230\pm 0.0007$
PBR (ours) $0.0630\pm 0.0045$ $0.8228\pm 0.0006$
PBR total (ours) $0.0618\pm 0.0054$ $0.8229\pm 0.0007$
PBR app. (ours) $0.0639\pm 0.0008$ $\mathbf{0.8230\pm 0.0007}$
PBR app. total (ours) $0.0639\pm 0.0008$ $0.8230\pm 0.0007$
Temperature scaling [10] $\mathbf{0.0222\pm 0.0013}$ $0.8230\pm 0.0007$